64r^2-9=0 Please solve by factoring. Thank you
Answer by ewatrrr(24785)

 
Hi,
64r^2-9=0
(8r - 3)(8r + 3) = 0 , r = ± 3/8
OR
64r^2-9=0
64r^2 = 9
r^2 = 9/64
r = ± 3/8